Define the following:

AD function with G & T

            Disposable income (Yd)

            Fiscal policy – expansionary & contractionary

            AD model impacts of fiscal policies

            Spending, Tax, and Balanced Budget Multipliers

            Automatic stabilizers

            Discretionary policy

            Deficit, Surplus (Budget)
Answer #1

AD function with G & T

Y = a+b(Y-T) +I+G , Increase in G and reduction T would cause AD to rise. or it shifts to right.

Disposable income (Yd)

Disposable income is income which left after paying off all personal taxes. it is difference between personal income and personal taxes. DI = Personal income - Personal taxes.

Fiscal policy – expansionary & contractionary

Fiscal policy means use of tax and expenditure to change direction of economy. Expansionary fiscal policy occurs when government raises expenditure and slash down taxes. Furthermore, contractionary policy means rise in taxes and fall in expenditure to ease down pressure of inflation.

AD model impacts of fiscal policies

Government seeks to influence AD through the taxes and expenditure. During the recession, AD shifts to left, hence government seeks to shift it to left.

Spending, Tax, and Balanced Budget Multipliers

Government needs to reduces its expenditure and raises taxes to achieve objective of balanced budget multiplier.

Tax multiplier = MPC/1-MPC

Expenditure multiplier = 1/1-MPC

Balanced budget Multiplier = - MPC/1-MPC + 1/1-MPC

= 1-MPC/1-MPC

= 1

Automatic stabilizers

Automatic stabilizers works without government actions or maneuvering to stabilize economy. Progressive taxes, unemployment allowances are examples of automatic stabilizers. Taxes increase when economy booms and automatically tax burden falls when economy enters into phase of recession.

Discretionary policy

Discretionary policy is deliberate actions by government to correct disequilibrium in economy.

Deficit, Surplus (Budget)

Deficit occurs when Expenditure outpaces revenue of government. While surplus implies that government expenditure is less than its revenue.